# DINOv2: Learning Robust Visual Features without Supervision

### v1

@@ Abstract
The recent breakthroughs in natural language processing for model pretraining on large quantities of data have opened the way for similar foundation models in computer vision. These models could greatly simplify the use of images in any system by producing all-purpose visual features, i.e., features that work across image distributions and tasks without finetuning. This work shows that existing pretraining methods, especially self-supervised methods, can produce such features if trained on enough curated data from diverse sources. We revisit existing approaches and combine different techniques to scale our pretraining in terms of data and model size. Most of the technical contributions aim at accelerating and stabilizing the training at scale. In terms of data, we propose an automatic pipeline to build a dedicated, diverse, and curated image dataset instead of uncurated data, as typically done in the self-supervised literature. In terms of models, we train a ViT model (Dosovitskiy et al., 2020) with 1B parameters and distill it into a series of smaller models that surpass the best available all-purpose features, OpenCLIP (Ilharco et al., 2021) on most of the benchmarks at image and pixel levels.
@@ 1 Introduction
Learning task-agnostic pretrained representations have become the standard in Natural Language Processing (NLP) (Radford et al., ; Raffel et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022; Touvron et al., 2023). One can use these features “as they are”, i.e., without fine-tuning, and achieve performances on downstream tasks that are significantly better than those produced by task-specific models (Brown et al., 2020). This success has been fueled by pretraining on large quantities of raw text using pretext objectives, such as language modeling (Radford et al., 2017) or word vectors (Devlin et al., 2018), that require no supervision.
@@
An alternative to text-guided pretraining is self-supervised learning  (Caron et al., 2018; Chen et al., 2020; He et al., 2021) where features are learned from images alone. These approaches are conceptually closer to pretext tasks such as language modeling and can capture information at the image and pixel level (Caron et al., 2021). However, despite their potential to learn all-purposed features, most of the advances in self-supervised learning were made in the context of pretraining on a small curated dataset, ImageNet-1k (Russakovsky et al., 2015). Some efforts on scaling these approaches beyond ImageNet-1k have been attempted (Caron et al., 2019; Goyal et al., 2021; 2022a), but they focused on uncurated datasets, which typically lead to a significant drop in the quality of the features. This is explained by the lack of control over the data quality and diversity, which are essential to produce good features.
@@
Regarding pretraining data, we have built an automatic pipeline to filter and rebalance datasets from an extensive collection of uncurated images. This pipeline is inspired by pipelines used in NLP (Wenzek et al., 2019), where data similarities are used instead of external metadata and do not require manual annotation. A major difficulty when dealing with images in the wild is to rebalance concepts and avoid overfitting on a few dominant modes. In this work, a naive clustering approach works reasonably well to resolve this issue. We gathered a small but diverse corpus of 142M images to validate our approach.
@@ 2 Related Work
A first family of self-supervised methods focuses on pretext tasks built from the image, i.e., extracting a signal from the image to be predicted from the rest of the image. This idea has become prevalent with the work of Doersch et al. (2015), where they train by predicting the context of a given patch. Many other pretext tasks were introduced based on re-colorizing images (Zhang et al., 2016), predicting transformations (Gidaris et al., 2018), inpainting (Pathak et al., 2016) or patch re-ordering (Noroozi & Favaro, 2016; Misra & Maaten, 2020). Recently, the emergence of patch-based architectures, like ViTs, has led to a revisit of inpainting for pre-training (He et al., 2021; Bao et al., 2021; El-Nouby et al., 2021), potentially in feature space (Assran et al., 2023; Baevski et al., 2022). Of particular interest, He et al. (2021) show that a masked auto-encoder (MAE) learns features that provide substantial improvements when finetuned on downstream tasks. This property of MAEs has been further validated on video (Tong et al., 2022), audio (Xu et al., 2022), and across other modalities (Girdhar et al., 2022). However, their features require supervised finetuning, while our features perform well out of the box.
@@
The second line of work, closer to ours, is using discriminative signals between images or groups of images to learn features. This family of methods has roots in early deep learning work (Hadsell et al., 2006) but became popular with the emergence of instance classification methods (Dosovitskiy et al., 2014; Bojanowski & Joulin, 2017; Wu et al., 2018). Several improvements were made based either on instance-level objectives (Hénaff et al., 2019; He et al., 2020; Chen & He, 2020; Chen et al., 2020; Grill et al., 2020; Caron et al., 2021) or clustering (Caron et al., 2018; Asano et al., 2020; Caron et al., 2020). These methods provide performant frozen features on standard benchmarks like ImageNet (Russakovsky et al., 2015), but they are hard to scale to larger model sizes (Chen et al., 2021). In this work, we revisit the training of these approaches in the context of large pretraining datasets and models. In particular, we build on top of Zhou et al. (2021) that we find particularly suited for scaling.
@@
Our dataset construction borrows from the image retrieval community (Weinzaepfel et al., 2021; Radenović et al., 2018b; Berman et al., 2019; Douze et al., 2009; Tolias et al., 2015; Revaud et al., 2019). In particular, the use of retrieval to augment the training set has been studied in the context of semi-supervised learning (Yalniz et al., 2019). Similarly, others have used hashtags or other metadata to filter uncurated datasets (Mahajan et al., 2018; Radford et al., 2021). Unlike this work, we use no metadata nor supervision to filter images and leverage visual similarity between images. Our approach is inspired by text curation pipelines (Wenzek et al., 2019), where a language model is trained on Wikipedia to score texts extracted from an uncurated source.
@@ 3 Data Processing
Data sources. Our selection of curated datasets is detailed in the appendix (Table 15) and contains ImageNet-22k, the train split of ImageNet-1k, Google Landmarks and several fine-grained datasets. For the uncurated data source, we collect a raw unfiltered dataset of images from a publicly available repository of crawled web data. From each web page in the repository, we extract URL links of images from <img> tags. We discards URLs that are unsafe or restricted by domains, and post-process the downloaded images (PCA hash deduplication, NSFW filtering, and blurring identifiable faces). This results in 1.2B unique images.
@@
Self-supervised image retrieval. We build our curated pretraining dataset by retrieving images from our uncurated data source that are close to images in our curated sources. In order to do this, we first compute an image embedding using a self-supervised ViT-H/16 network pretrained on ImageNet-22k, and use cosine-similarity as a distance measure between images. Then, we perform k-means clustering of the uncurated data. Given a query dataset for retrieval, if it is large enough we retrieve $N$ (typically 4) nearest neighbors for each query image. If it is small, we sample $M$ images from the cluster corresponding to each query image. We adjust $N$ and $M$ by visual inspection of the retrieval result.
@@ 4 Discriminative Self-supervised Pre-training
  • Image-level objective (Caron et al., 2021). We consider the cross-entropy loss between the features extracted from a student and a teacher network. Both features are coming from the class token of a ViT, obtained from different crops of the same image. We learn the parameters of the student and build the teacher with an exponential moving average of past iterates (He et al., 2020).
  • Patch-level objective (Zhou et al., 2021). We randomly mask some of the input patches given to the student, but not to the teacher. We then add a cross-entropy loss between the patch features of both networks on each masked patch. This loss is combined with the image-level loss.
  • Untying head weights between both objectives. We observe that tying the weights associated with both objectives makes the model underfit at the patch-level while overfitting at the image-level. Untying these weights resolves this issue and improve the performances at both scales.
@@
  • KoLeo regularizer  (Sablayrolles et al., 2018). The KoLeo regularizer derives from the Kozachenko-Leonenko differential entropy estimator (see  Beirlant et al. (1997); Delattre & Fournier (2017)) and encourages a uniform span of the features within a batch. Given a set of $n$ vectors $(x_{1},\dots,x_{n})$, it is defined as ${\mathcal{L}}_{\mathrm{koleo}}=-\frac{1}{n}\sum_{i=1}^{n}\log(d_{n,i}),$ where $d_{n,i}=\min_{j\neq i}\|x_{i}-x_{j}\|$ is the minimum distance between $x_{i}$ and any other point within the batch. We also $\ell_{2}$-normalize the features before computing this regularizer.
  • Adapting the resolution (Touvron et al., 2019). Increasing image resolution is key to pixel-level downstream tasks such as segmentation or detection, where small objects disappear at low resolutions. However, training at high resolution is time and memory demanding, and instead, we increase the resolution of images to $518\times 518$ during a short period at the end of pretraining.
@@ 5 Efficient implementation
We consider several improvements to train models at a larger scale. We train models on A100 GPUs using PyTorch 2.0. The code is also available along with the pretrained models used for feature extraction  [^1] https://github.com/facebookresearch/dinov2. The details of our models are in the appendix, Table 17. With the same hardware, compared to the iBOT implementation, the DINOv2 code runs around $2\times$ faster using only $1/3$ of the memory.
@@
Nested tensors in self-attention.
Our version also allows running in the same forward pass the global crops and the local crops (that have different numbers of patch tokens), leading to significant compute efficiency gains compared to using separate forward and backward passes as done in prior implementations. The lower-level components of our setup are available in the xFormers library [^2] https://github.com/facebookresearch/xformers (Lefaudeux et al. (2022)).
@@
Minimizing our objective with the AdamW optimizer requires 4 model replicas in float32 precision – student, teacher, optimizer first moments, optimizer second moments. This sums to $16~\mathrm{GB}$ of memory for a billion-parameter model such as our ViT-g. In order to reduce this memory footprint per GPU, we split the model replicas across GPUs, i.e., sharding $16~\mathrm{GB}$ across GPUs using the PyTorch implementation of FSDP. Consequently, the model size is not bounded by the memory of a single GPU but by the total sum of GPU memory across compute nodes. The Pytorch implementation of FSDP brings a second advantage, which is to save on the cross-GPU communication costs: the weight shards are stored in float32 precision as required by the optimizer, but broadcasting weights and reducing gradients is done in float16 precision for the backbone (MLP heads gradients are reduced in float32 to avoid training instabilities). This leads to approximately 50% reduction in communication costs compared to the float32 gradient all-reduce operation used in DistributedDataParallel (DDP), which is used in other self-supervised pretraining methods (Caron et al., 2021; Zhou et al., 2021). As a consequence, the training procedure scales more efficiently than DDP with float16 autocast when scaling the number of GPU nodes. Overall, Pytorch-FSDP mixed-precision is superior to DDP with autocast in virtually all cases we encountered.
@@
Most of our technical improvements to the training loop aim at improving the training of large models over large quantities of data. For smaller models, we distill them from our largest model, the ViT-g, instead of training them from scratch. Knowledge distillation (Hinton et al., 2015) aims at reproducing the output of a large model with a smaller model by minimizing some distance between both outputs for a set of given inputs. Since our objective function is a form of distillation from the teacher network to the student network, we leverage the same training loop with a few exceptions: we use a larger model as a frozen teacher, keep a spare EMA of the student that we use as our final model, remove the masking and stochastic depth, and, apply the iBOT loss on the two global crops. In our ablations, we observe that this approach achieves better performance than training from scratch, even for a ViT-L. Our distillation method ends up close to the one described by Duval et al. (2023), except we do not modify the loss terms for distillation and evaluate the EMA of the student.
@@ 6.1 Improved Training Recipe
Our approach improves over the iBOT method by combining it with several existing components described in Sec. 4. To evaluate their importance, we train multiple models where we successively add components to a baseline iBOT model. We report the Top-1 accuracy on the validation set of ImageNet-1k with a k-NN and a linear linear in Table 1. Generally, we observe that each component improves the performance on either k-NN or linear probing and even both in most cases. Only LayerScale and Stochastic Depth incur a performance drop in linear probing but significantly improve the training stability in our experience.
@@
Table 1: Ablation study of the training differences between iBOT and DINOv2. We optimize for k-NN performance, as in our experience, the linear probe performance is lower-bounded by the k-NN performance. Some modifications, like LayerScale and a high Stochastic Depth (rate=$0.4$), incur a decrease in linear probe performance, but have the benefits of increasing the stability of training by avoiding NaN loss values during training. Overall, these modifications allowed for the next set of improvements to be added. Experiments are run using the ViT-Large architecture on ImageNet-22k.

### v2

@@ Abstract
The recent breakthroughs in natural language processing for model pretraining on large quantities of data have opened the way for similar foundation models in computer vision. These models could greatly simplify the use of images in any system by producing general-purpose visual features, i.e., features that work across image distributions and tasks without finetuning. This work shows that existing pretraining methods, especially self-supervised methods, can produce such features if trained on enough curated data from diverse sources. We revisit existing approaches and combine different techniques to scale our pretraining in terms of data and model size. Most of the technical contributions aim at accelerating and stabilizing the training at scale. In terms of data, we propose an automatic pipeline to build a dedicated, diverse, and curated image dataset instead of uncurated data, as typically done in the self-supervised literature. In terms of models, we train a ViT model (Dosovitskiy et al., 2021) with 1B parameters and distill it into a series of smaller models that surpass the best available general-purpose features, OpenCLIP (Ilharco et al., 2021) on most of the benchmarks at image and pixel levels.
@@ 1 Introduction
Learning task-agnostic pretrained representations have become the standard in Natural Language Processing (NLP) (Radford et al., 2019; Raffel et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022; Touvron et al., 2023). One can use these features “as they are”, i.e., without fine-tuning, and achieve performances on downstream tasks that are significantly better than those produced by task-specific models (Brown et al., 2020). This success has been fueled by pretraining on large quantities of raw text using pretext objectives, such as language modeling (Radford et al., 2017) or word vectors (Devlin et al., 2019), that require no supervision.
@@
An alternative to text-guided pretraining is self-supervised learning  (Caron et al., 2018; Chen et al., 2020; He et al., 2022) where features are learned from images alone. These approaches are conceptually closer to pretext tasks such as language modeling and can capture information at the image and pixel level (Caron et al., 2021). Additionally, the features output by self-supervised models have been shown to exhibit various useful properties, and have enabled enabled a wide variety of applications (Amir et al., 2022; Tumanyan et al., 2022; Ofri-Amar et al., 2023; Hamilton et al., 2022). However, despite their potential to learn general-purpose features, most of the advances in self-supervised learning were made in the context of pretraining on a small curated dataset, ImageNet-1k (Russakovsky et al., 2015). Some efforts on scaling these approaches beyond ImageNet-1k have been attempted (Caron et al., 2019; Goyal et al., 2021; 2022a), but they focused on uncurated datasets, which typically lead to a significant drop in the quality of the features. This is explained by the lack of control over the data quality and diversity, which are essential to produce good features.
@@
Regarding pretraining data, we have built an automatic pipeline to filter and rebalance datasets from an extensive collection of uncurated images. This pipeline is inspired by pipelines used in NLP (Wenzek et al., 2020), where data similarities are used instead of external metadata and do not require manual annotation. A major difficulty when dealing with images in the wild is to rebalance concepts and avoid overfitting on a few dominant modes. In this work, a naive clustering approach works reasonably well to resolve this issue. We gathered a small but diverse corpus of 142M images to validate our approach.
@@ 2 Related Work
A first family of self-supervised methods focuses on pretext tasks built from the image, i.e., extracting a signal from the image to be predicted from the rest of the image. This idea has become prevalent with the work of Doersch et al. (2015), where they train by predicting the context of a given patch. Many other pretext tasks were introduced based on, for example, re-colorizing images (Zhang et al., 2016), predicting transformations (Gidaris et al., 2018), inpainting (Pathak et al., 2016) or patch re-ordering (Noroozi & Favaro, 2016; Misra & Maaten, 2020). Recently, the emergence of patch-based architectures, like ViTs, has led to a revisit of inpainting for pre-training (He et al., 2022; Bao et al., 2021; El-Nouby et al., 2021), potentially in feature space (Assran et al., 2023; Baevski et al., 2022). Of particular interest, He et al. (2022) show that a masked auto-encoder (MAE) learns features that provide substantial improvements when finetuned on downstream tasks. This property of MAEs has been further validated on video (Tong et al., 2022), audio (Xu et al., 2022), and across other modalities (Girdhar et al., 2023). However, their features require supervised finetuning, while our features perform well out of the box.
@@
The second line of work, closer to ours, is using discriminative signals between images or groups of images to learn features. This family of methods has roots in early deep learning work (Hadsell et al., 2006) but became popular with the emergence of instance classification methods (Dosovitskiy et al., 2016; Bojanowski & Joulin, 2017; Wu et al., 2018). Several improvements were made based either on instance-level objectives (Hénaff et al., 2019; He et al., 2020; Chen & He, 2021; Chen et al., 2020; Grill et al., 2020; Caron et al., 2021) or clustering (Caron et al., 2018; Asano et al., 2020; Caron et al., 2020). These methods provide performant frozen features on standard benchmarks like ImageNet (Russakovsky et al., 2015), but they are hard to scale to larger model sizes (Chen et al., 2021). In this work, we revisit the training of these approaches in the context of large pretraining datasets and models. In particular, we build on top of Zhou et al. (2022a) that we find particularly suited for scaling.
@@
Our dataset construction borrows from the image retrieval community (Weinzaepfel et al., 2021; Radenović et al., 2018b; Berman et al., 2019; Douze et al., 2009; Tolias et al., 2016; Revaud et al., 2019). In particular, the use of retrieval to augment the training set has been studied in the context of semi-supervised learning (Yalniz et al., 2019). Similarly, others have used hashtags or other metadata (Mahajan et al., 2018; Radford et al., 2021) or pretrained vision encoders (Schuhmann et al., 2021; 2022) to filter uncurated datasets. Unlike these works, we use no pretrained encoders, metadata nor supervision to filter images and leverage visual similarity between images. Our approach is inspired by text curation pipelines (Wenzek et al., 2020), where a language model is trained on Wikipedia to score texts extracted from an uncurated source.
@@ 3 Data Processing
Data sources. Our selection of curated datasets is detailed in the appendix (Table 15) and contains ImageNet-22k, the train split of ImageNet-1k, Google Landmarks and several fine-grained datasets. For the uncurated data source, we collect a raw unfiltered dataset of images from a publicly available repository of crawled web data. From each web page in the repository, we extract URL links of images from <img> tags. We discard URLs that are unsafe or restricted by domains, and post-process the downloaded images (PCA hash deduplication, NSFW filtering, and blurring identifiable faces). This results in 1.2B unique images.
@@
Self-supervised image retrieval. We build our curated pretraining dataset by retrieving images from our uncurated data source that are close to images in our curated sources. In order to do this, we first compute an image embedding using a self-supervised ViT-H/16 network pretrained on ImageNet-22k, and use cosine-similarity as a distance measure between images. Then, we perform k-means clustering of the uncurated data. Given a query dataset for retrieval, if it is large enough we retrieve $N$ (typically 4) nearest neighbors for each query image. If it is small, we sample $M$ images from the cluster corresponding to each query image. Although visual inspection seemed to indicate good retrieval quality for $N$ much larger than 4, this leads to more collisions (images that are nearest-neighbor retrievals of multiple queries). We choose $N=4$ as it provides a good tradeoff in that sense.
@@ 4 Discriminative Self-supervised Pre-training
  • Image-level objective (Caron et al., 2021). We consider the cross-entropy loss between the features extracted from a student and a teacher network. Both features are coming from the class token of a ViT, obtained from different crops of the same image. We pass the student class token through the student DINO head. This head is an MLP model outputting a vector of scores, that we call "prototype scores". We then apply a softmax to obtain $p_{s}$. Similarly, we apply the teacher DINO head to the teacher class token to obtain teacher prototype scores. We then apply a softmax followed by a centering with moving average (or a Sinkhorn-Knopp centering as detailed thereafter) to obtain $p_{t}$. The DINO loss term corresponds to:
    ${\mathcal{L}}_{DINO}=-\sum p_{t}\log p_{s}$
    We learn the parameters of the student and build the teacher head with an exponential moving average of past iterates (He et al., 2020).
  • Patch-level objective (Zhou et al., 2022a). We randomly mask some of the input patches given to the student, but not to the teacher. We then apply the student iBOT head to the student mask tokens. Similarly, we apply the teacher iBOT head to the (visible) teacher patch tokens corresponding to the ones masked in the student. We then apply the softmax and centering steps as above, and obtain the iBOT loss term:
    ${\mathcal{L}}_{iBOT}=-\sum_{i}p_{ti}\log p_{si}$
    , where $i$ are patch indices for masked tokens. Similarly to above, we learn the parameters of the student, and build the teacher head through exponential moving average.
  • Untying head weights between both objectives. Both the DINO and the iBOT loss use a learnable MLP projection head. It is applied to the output tokens and the loss is compute atop. In Zhou et al. (2022a), an ablation study shows that sharing parameters between the DINO and iBOT heads leads to better performance. At scale, we observed that the opposite is true, and we therefore use two separate heads in all our experiments.
@@
  • KoLeo regularizer  (Sablayrolles et al., 2019). The KoLeo regularizer derives from the Kozachenko-Leonenko differential entropy estimator (see  Beirlant et al. (1997); Delattre & Fournier (2017)) and encourages a uniform span of the features within a batch. Given a set of $n$ vectors $(x_{1},\dots,x_{n})$, it is defined as
    ${\mathcal{L}}_{\mathrm{koleo}}=-\frac{1}{n}\sum_{i=1}^{n}\log(d_{n,i}),$
    where $d_{n,i}=\min_{j\neq i}\|x_{i}-x_{j}\|$ is the minimum distance between $x_{i}$ and any other point within the batch. We also $\ell_{2}$-normalize the features before computing this regularizer.
  • Adapting the resolution (Touvron et al., 2019). Increasing image resolution is key to pixel-level downstream tasks such as segmentation or detection, where small objects disappear at low resolutions. However, training at high resolution is time and memory demanding, and instead, we increase the resolution of images to $518\times 518$ during a short period at the end of pretraining. This is also similar to UniViT training from Likhomanenko et al. (2021) and FlexiViT training from Beyer et al. (2023).
@@ 5 Efficient implementation
We consider several improvements to train models at a larger scale. We train models on A100 GPUs using PyTorch 2.0. The code and pretrained models are made available under Apache 2.0 license  [^1] https://github.com/facebookresearch/dinov2. The details of our models are in the appendix, Table 17. With the same hardware, compared to the iBOT implementation, the DINOv2 code runs around $2\times$ faster using only $1/3$ of the memory.
@@
Sequence packing.
The DINO algorithm requires forwarding both large crops (at resolution 224) and small crops (resolution 98). When split into patches, these two groups are represented by token sequences of different lengths and cannot be forwarded together. In order to accelerate training, we use a trick called "sequence packing," which originates from NLP (Krell et al., 2022). The idea is simple: we concatenate the sequences we must forward through the transformers into a single long sequence. We pass this sequence through the transformer blocks as usual. However, a block-diagonal mask is applied to the self-attention matrix in attention layers, preventing attention between different sequences. This way, the forward is strictly equivalent to forwarding each sequence separately. This trick gives us significant compute efficiency gains compared to using separate forward and backward passes, as in prior implementations. The lower-level components of our setup are available in the xFormers library [^2] https://github.com/facebookresearch/xformers (Lefaudeux et al. (2022)).
@@
Minimizing our objective with the AdamW optimizer requires 4 model replicas in float32 precision – student, teacher, optimizer first moments, optimizer second moments. This sums to $16~\mathrm{GB}$ of memory for a billion-parameter model such as our ViT-g. In order to reduce this memory footprint per GPU, we split the model replicas across GPUs, i.e., sharding $16~\mathrm{GB}$ across GPUs using the PyTorch implementation of FSDP. Consequently, the model size is not bounded by the memory of a single GPU but by the total sum of GPU memory across compute nodes. The Pytorch implementation of FSDP brings a second advantage, which is to save on the cross-GPU communication costs: the weight shards are stored in float32 precision as required by the optimizer, but broadcasting weights and reducing gradients is done in float16 precision for the backbone (MLP heads gradients are reduced in float32 to avoid training instabilities). This leads to approximately 50% reduction in communication costs compared to the float32 gradient all-reduce operation used in DistributedDataParallel (DDP), which is used in other self-supervised pretraining methods (Caron et al., 2021; Zhou et al., 2022a). As a consequence, the training procedure scales more efficiently than DDP with float16 autocast when scaling the number of GPU nodes. Overall, Pytorch-FSDP mixed-precision is superior to DDP with autocast in virtually all cases we encountered.
@@
Most of our technical improvements to the training loop aim at improving the training of large models over large quantities of data. For smaller models, we distill them from our largest model, the ViT-g, instead of training them from scratch. Knowledge distillation (Hinton et al., 2014) aims at reproducing the output of a large model with a smaller model by minimizing some distance between both outputs for a set of given inputs. Since our objective function is a form of distillation from the teacher network to the student network, we leverage the same training loop with a few exceptions: we use a larger model as a frozen teacher, keep a spare EMA of the student that we use as our final model, remove the masking and stochastic depth, and, apply the iBOT loss on the two global crops. In our ablations, we observe that this approach achieves better performance than training from scratch, even for a ViT-L. Our distillation method ends up close to the one described by Duval et al. (2023), except we do not modify the loss terms for distillation and evaluate the EMA of the student.
@@ 6.1 Improved Training Recipe
Our approach improves over the iBOT method by combining it with several existing components described in Sec. 4. To evaluate their importance, we train multiple models where we successively add components to a baseline iBOT model. We report the Top-1 accuracy on the validation set of ImageNet-1k with a k-NN and a linear probe in Table 1. Generally, we observe that each component improves the performance on either k-NN or linear probing and even both in most cases. Only LayerScale and Stochastic Depth incur a performance drop in linear probing but significantly improve the training stability in our experience.
@@
Table 1: Ablation study of the training differences between iBOT and DINOv2. We optimize for k-NN performance, as in our experience, the linear probe performance is lower-bounded by the k-NN performance. Some modifications, like LayerScale and a high Stochastic Depth (rate=$0.4$), incur a decrease in linear probe performance, but have the benefits of increasing the stability of training by avoiding NaN loss values during training (Touvron et al., 2022). Overall, these modifications allowed for the next set of improvements to be added. Experiments are run using the ViT-Large architecture on ImageNet-22k.